OpenStudy (anonymous):

True of False 1.The equation of every line can be written in point-slope form. 2. If two nonvertical lines are parallel, they have the same slope 3.If the x and y intercepts of a line ar erational and nonzero, then the slope of the line is rational 4.(3x-2y+4)+m(2x+6y-2) = 0 is the equation of a line for each real number m. 5.The natural domain of T(theta) = sec(theta) + cos(theta) 6.The range of the function f(x) = tanx - secx is the set (-inf,-1]U[1,inf).

OpenStudy (raffle_snaffle):

Well what do you think?

OpenStudy (tkhunny):

Please provide counter examples to all those things you consider false.

OpenStudy (anonymous):

1. I believe it is true. 2. True, since that is what makes them parallel 3.Honestly I just don't understand the question there. I would suppose it is rational? 4.Again, don't really see what it is saying. 5.Looking at the graph I want to say True. 6.Looking at the graph, I think false.

OpenStudy (tkhunny):

Nice Try. Consider a vertical line and try #1 again.

OpenStudy (tkhunny):

What has the intercept to do with the slope? Nothing! Forget #3 FALSE.

OpenStudy (anonymous):

so y = x. I guess that wouldn't be in that form.

OpenStudy (tkhunny):

No, y = x is in Slope-Intercept Form. x = 2 is not, and cannot be in Slope-Intercept form.

OpenStudy (triciaal):

@tkhunny are you saying all lines cannot be written in point-slope format?

OpenStudy (tkhunny):

When did I say that? I said no VERTICAL line can be written in Slope-Intercept from.

OpenStudy (aum):

#1 is false due to the fact that for vertical lines, slope is undefined.

OpenStudy (anonymous):

I think I got 1-3 down..still stumped on 4-6..

OpenStudy (tkhunny):

#4 just wants to know if there is a value for 'm' that will make that expression NOT an equation of a line. (3x-2y+4)+m(2x+6y-2) = 0 Let's try Slope-Intercept Form. It will be fun. (3x-2y+4)+m(2x+6y-2) = 0 3x + 2mx - 2y + 6my + 4 - 2m = 0 6my - 2y = -3x - 2mx - 4 + 2m y(6m-2) = x(-3-2m) + (2m-4) y = x(-3-2m)/(6m-2) + (2m-4)/(6m-2) Wasn't that as fun as I said? The question is almost, is "Can 6m-2 = 0"? That would be no good. But would that just be a vertical line?

OpenStudy (aum):

what was your answer for 3?

OpenStudy (tkhunny):

What's \(T(\theta)\)?

OpenStudy (tkhunny):

For #6, you may wish to convert to sine and cosine and get a different look at it.

OpenStudy (anonymous):

That's all it says. The natural domain of T(Theta) = sec(theta) + cos(theta) is (-inf,inf)

OpenStudy (anonymous):

For number 3 I have false.

OpenStudy (aum):

For #3 it is false most of the times but there may be exceptions. If the y-intercept is \(\sqrt{32}\) and the x-intercept is \(\sqrt{2}\), couldn't the slope be rational?

OpenStudy (aum):

For #5, the domain for T is same as the domain for the secant function.

OpenStudy (tkhunny):

#3 We only get TRUE or FALSE. There is no CONDITIONAL. It may be that it is so, but there is no inherent relationship. It is FALSE. #5, I see. A "Natural Domain" is the Domain that has no arbitrary restriction. \(T(\theta) = \sec(\theta) + \cos(theta) = \dfrac{1}{\cos(\theta)} + \cos(\theta)\) It does appear that we are missing some things. What happens when \(\cos(\theta) = 0\)?

OpenStudy (anonymous):

I think I'll stick with false for it. I mean, I'm not getting confirmation either way, but I think it is false as well.

OpenStudy (aum):

Plotting #6 shows the range as \((-\infty, \infty)\).

OpenStudy (anonymous):

So that one would be false? I'm getting really frustrated at these since I try working them and the homework thing gives me a wrong answer and doesn't specify which is wrong. Feel like I'm just feeling around in the dark here.

OpenStudy (anonymous):

So far, the ones I feel like I KNOW are right are 1.F, 2.T, 3.F,6.F

OpenStudy (aum):

Try 5F and 4T (there are exceptions to 4 but it is mostly true).

OpenStudy (anonymous):

No dice.

OpenStudy (aum):

How many tries are you allowed? Try 5F and 4F.

OpenStudy (anonymous):

Unlimited. I've already attempted 40 =P

OpenStudy (anonymous):

Still nothing D: . Feels like I've tried every combination up to this point, haha.

OpenStudy (aum):

Just for fun try 3F.

OpenStudy (aum):

That is 3F, 4F, 5F.

OpenStudy (anonymous):

I already had 3 as F I think. but still, no go.

OpenStudy (aum):

There are only 64 possibilities for all 6 questions. But a few of the questions are unambiguous and their answers can be fixed. And for every answer that is known for sure, the number possibilities will be lowered by a factor of 2.

OpenStudy (aum):

gtg.

OpenStudy (anonymous):

Holy crap I figured it! I just started messing around and got it.

OpenStudy (anonymous):

1.F 2.T 3.T 4.T 5.F 6.F

OpenStudy (anonymous):

Thanks for sticking around as long as you did!

OpenStudy (aum):

You are welcome. Nice of you to post the final answers although 3 is a bit controversial. I gave an example of an exception earlier that would make it true but there are many for which it is false.